mariah16:

6 (-2 + 5x) - 6x = -180

mariah16:

i got the first part then i got lost ;-;

mariah16:

so what i got so far is 12-+30

cookieoreo11:

this is so ๐Ÿ˜ƒ

cookieoreo11:

confusing

mariah16:

6(-2+5x)-6x=-180

supie:

you would start of by solving \(6(-2+5x)\) you can solve that by using the distributive property, the distributive property is when you multiply the number that is outside of the parenthesis by everything that is inside of the parenthesis, \(a(b+c)=ab+ac\) so, using the distributive property to solve \(6(-2+5x)\) you would be multiplying 6 by the numbers that are inside of the parenthesis, \(6\times-2 =?\\6\times5x=?\) do you think you can solve that?

cookieoreo11:

thank you

cookieoreo11:

i think i can

cookieoreo11:

confused btw i have the same work as her

mariah16:

you still confused?

cookieoreo11:

yes

mariah16:

so 6 x -2 = -12

cookieoreo11:

OHHHh

mariah16:

and 6 x 5 = 30

joshdeanda16:

who knows what -4(-3+6x)-4x =180 who knows the answer for this

mariah16:

so its -12 + 30x

supie:

@mariah16 wrote:
and 6 x 5 = 30
it's not \(6\times5\) it's \(6\times5\color{red}{x}=30x\)

cookieoreo11:

ayo supie thank you fot helping

supie:

@mariah16 wrote:
so its -12 + 30x
yeah good job e.e

cookieoreo11:

for*

supie:

@cookieoreo11 wrote:
ayo supie thank you fot helping
you're welcome yk you're not done tho, right?

cookieoreo11:

yeah

supie:

alr so you still need help or no- ?

cookieoreo11:

yes

supie:

@mariah16 wrote:
so its -12 + 30x
ok well as mariah16 said, you get -12+30x so we're left with \(โˆ’12+30x+(โˆ’6x)=โˆ’180\) so then you would combine the like terms, which are 30x and -6x, \(30x+(-6x)\) you think you can solve 30x+(-6x)?

cookieoreo11:

yes

supie:

k then do it lol

cookieoreo11:

36? or - 36?

supie:

no, remember that you're adding negative 6x not just 6x so you're technically subtracting 30x-6x

cookieoreo11:

ohhh okay

supie:

so 30x-6x= what

cookieoreo11:

24

cookieoreo11:

or

cookieoreo11:

-24

supie:

both are technically wrong, it would actually be 24x because 30-6=24 so 30x-6x=24x you did good but you just forgot the x so now you're left with the equation \(24xโˆ’12=โˆ’180\) what you would do is get `-12` and change the minus to a plus then add 12 to both sides of the equation you would be doing \(24x\cancel{\color{Red}{โˆ’12+12}}=\color{aqua}{โˆ’180+12}\) -12+12 cancels, what's -180+12?

cookieoreo11:

6x

cookieoreo11:

or - 6 x

supie:

sorry i was afk and it's -180+12 not -18+12

supie:

but it would give you \(โˆ’168\) and we're left with \(24x=โˆ’168\)

supie:

the next step is to divide both sides of \(24x=โˆ’168\) by 24 \(\large\cancel{\frac{24x}{24}}=\frac{-168}{24}\) 24x/2x cancels, because 24รท24=1 and we're left with 1x, and 1x=x so we have \(\large x=\frac{-168}{24}\) and \(-160รท24=-7\) so, \(x=-7\)